OpenStudy (anonymous):

Mike and Jamal are 9 miles apart, and are planning to meet up. Mike is walking at an average speed of 3 miles per hour to meet Jamal. Jamal is driving at an average speed of 25 miles per hour to meet Mike.

OpenStudy (md152727):

what is the question?

OpenStudy (anonymous):

Which equation can be used to find t, the time it takes for Mike and Jamal to meet? 25t – 3t = 0 25t – 3t = 9 25t + 3t = 1 25t + 3t = 9

OpenStudy (anonymous):

\[25 t+3 t=9 \] t = d/r t = 9/(3+25)=19.29 minutes